# SAQR-QC: A Logic for Scalable but Approximate Quantitative Reasoning about Quantum Circuits

### v1

@@ 5. Quantitative Reasoning of general GHZ circuit
This section analyzes a generalized GHZ circuit in which half of the gates are arbitrary single-qubit unitaries, resulting in a highly entangled and densely parameterized system that lies beyond the reach of the Gottesman–Knill theorem (Aaronson_2004, ). We evaluate several quantitative properties of the circuit’s output state using the techniques developed in this paper. In particular, we apply the “warm-up exercise” method from § 4.7 (Equation 17) to support scalable reasoning throughout the analysis. In the final step, we invoke Quantum Abstract Interpretation (QAI)—in particular, make use of 3.2—to infer that the output state resides within a two-dimensional subspace. We then apply our quantitative-reasoning framework to approximate the amplitudes of the basis vectors spanning this subspace. Notably, the resulting characterization of the output state is accurate up to a global phase factor when compared to the exact output vector. This example illustrates the point made near the end of § 1 that using SAQR-QC is not “cookbook.” The reason we obtain such a precise characterization of the output state is because of a careful choice of predicates.
Figure 3. GHZ circuit with one-qubit unitaries $U_{i}$.
We select the following domain ($(\{1,2\},\{2,3\},\cdots,\{n-1,n\})$.
STEP 1. We first select the precondition to be $\{\mathscr{A}|\mathcal{P}\}$ where
@@
After the first $H$ gate, we can use inequality 16 to derive the following:
$\sum_{i}{A_{s_{i}}}\otimes I_{[n]\setminus s_{i}}\leq H^{{\dagger}}\sum_{i}{B_{s_{i}}}\otimes I_{[n]\setminus s_{i}}H$
$\Longleftrightarrow$ $H\sum_{i}{A_{s_{i}}}\otimes I_{[n]\setminus s_{i}}H^{{\dagger}}\leq\sum_{i}{B_{s_{i}}}\otimes I_{[n]\setminus s_{i}}$
$\Longleftrightarrow$ $\sum_{i}H{A_{s_{i}}}\otimes I_{[n]\setminus s_{i}}H^{{\dagger}}\leq\sum_{i}{B_{s_{i}}}\otimes I_{[n]\setminus s_{i}}$
@@
The final inequality above matches inequality 18 on a term-by-term basis. We can satisfy inequality 18 when each of the following single-term inequalities hold:
@@
In this case, each of the inequalities can be satisfied as an equality, yielding the postcondition $\{\mathscr{B}|\mathcal{Q}\}$
@@
$\Longleftrightarrow$ $\sum_{i}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}\leq\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1}$
@@
$\Longleftrightarrow$ $\sum_{i\leq 2}\text{CNOT}_{1,2}(I_{[n]\setminus\{i,i+1\}}\otimes B_{i,i+1})\text{CNOT}_{1,2}^{{\dagger}}+\sum_{i>2}(\text{CNOT}_{1,2}I_{[n]\setminus\{i,i+1\}}\text{CNOT}_{1,2}^{{\dagger}}\otimes B_{i,i+1})\leq\sum_{i}I_{[n]\setminus\{i,i+1\}}\otimes C_{i,i+1}$
@@
At this point, we can satisfy inequality 18 by satisfying the following smaller inequalities:
@@
[…]
@@
(which happens to be the same as the pre-state $\mathscr{B}$ predicate $(B_{1,2},B_{2,3},\cdots,B_{n-1,n})$), and the post-state $\mathcal{R}$ predicate is
@@
Equation 22 illustrates an advantage of our choice of predicates. Because of properties such as those given in Equation 21, Equation 16 remains invariant under the application of CNOT gates. This invariance allows us to derive the strongest postcondition, while preserving the local structure of the matrix representation of predicates. As a result, we were able to make choices that made the inequalities that we worked with tight (or saturated, i.e., satified as equalities), making it easier to determine the postcondition.
The right-hand side of Equation 22 continues to serve as the predicate of local observables, as reasoning continues about the remaining CNOT gates.
After applying the CNOT gate on $q_{1},q_{r}$ for some $r$, we can choose the postcondition to be $(\mathscr{D}|\mathcal{S})$ where
[…]
@@
|$\ket{\beta_{1}}$ $=U_{1}\ket{0},\ \ \ket{\beta_{i}}=U_{i}\ket{+}\ \ \forall i>1$
|$\ket{\psi_{i}}$ $=U_{i}\ket{0},\ \ \ \ket{\phi_{i}}=U_{i}\ket{1}\ \ \forall i\geq 1.$
Let us denote the output state as $\rho=|\Psi\rangle\langle\Psi|$. According to 4.7, our proof of $\mathscr{F}$ implies that
$\sum_{i=1}^{n-1}{\mathrm{Tr}}(|00\rangle\langle 00||++\rangle\langle++|)\leq\sum_{i=1}^{n-1}{\mathrm{Tr}}[\rho_{i,i+1}(\beta_{i}\otimes\beta_{i+1})].$
That is
(23) $\frac{n-1}{4}\leq\sum_{i=1}^{n-1}{\mathrm{Tr}}[\rho_{i,i+1}(\beta_{i}\otimes\beta_{i+1})].$
STEP 2. This step is similar to Step 1, except that we start with the precondition
$((|--\rangle\langle--|,|--\rangle\langle--|,\cdots,|--\rangle\langle--|)|\mathcal{P})$. We can compute a postcondition of the whole circuit as
$((\delta_{1}\otimes\delta_{2}\cdots,{\delta_{n-1}\otimes\delta_{n}})|\mathcal{T}),$
where $\mathcal{T}$ is the same as in Step 2, and $\delta_{i}=|\delta_{i}\rangle\langle\delta_{i}|,\ \ \ket{\delta_{1}}=U_{1}\ket{1},\ \ \ket{\delta_{i}}=U_{i}\ket{-}\ \ \forall i>1$. These conditions imply that
@@
${\mathrm{Tr}}[(\psi_{1}\otimes\psi_{2})(\beta_{1}\otimes\beta_{2})]=\frac{1}{2},\ {\mathrm{Tr}}[(\psi_{i}\otimes\psi_{i+1})(\beta_{i}\otimes\beta_{i+1})]=\frac{1}{4}\ \ \forall 1<i$
${\mathrm{Tr}}[(\phi_{1}\otimes\phi_{2})(\beta_{1}\otimes\beta_{2})]=0,\ {\mathrm{Tr}}[(\phi_{i}\otimes\phi_{i+1})(\beta_{i}\otimes\beta_{i+1})]=\frac{1}{4}\ \ \forall 1<i$
${\mathrm{Tr}}[(\psi_{1}\otimes\psi_{2})(\delta_{1}\otimes\delta_{2})]=0,\ {\mathrm{Tr}}[(\psi_{i}\otimes\psi_{i+1})(\delta_{i}\otimes\delta_{i+1})]=\frac{1}{4}\ \ \forall 1<i$
${\mathrm{Tr}}[(\phi_{1}\otimes\phi_{2})(\delta_{1}\otimes\delta_{2})]=\frac{1}{2},\ {\mathrm{Tr}}[(\phi_{i}\otimes\phi_{i+1})(\delta_{i}\otimes\delta_{i+1})]=\frac{1}{4}\ \ \forall 1<i.$
Equations 23 and 24 imply
@@
(25) |$\ket{\Psi}=\frac{1}{\sqrt{2}}(\ket{\psi_{1}\cdots\psi_{n}}+e^{i\theta}\ket{\phi_{1}\cdots\phi_{n}}).$
Equation 25 represents a closed-form expression for the circuit’s output with a single unknown real parameter. Such quantitative reasoning cannot be achieved with qualitative methods like QAI alone.
@@ 6. Quantum Phase Estimation
This section applies our proof system to analyze Quantum Phase Estimation (QPE), a foundational algorithm in quantum computing. QPE plays a pivotal role in applications such as Shor’s algorithm, quantum simulation, and other tasks requiring precise eigenvalue estimation—serving as a key source of quantum speedup over classical approaches.
We begin by verifying the correctness of the Quantum Fourier Transform (QFT) using Quantum Abstract Interpretation (QAI). Building upon the QAI results, we then apply our logic framework to reason about the Quantum Phase Estimation (QPE) circuit. In particular, we use our proof system to derive a lower bound on the success probability of QPE, demonstrating the power of our scalable quantitative reasoning approach.
To the best of our knowledge, QFT and QPE have not previously been analyzed using any scalable formal-reasoning framework. Our method provides a novel and tractable approach to reasoning about the correctness and quantitative behavior of QPE circuits.
@@ 6.1. Quantum Fourier transform
The QFT is the quantum analog of the discrete Fourier transform, central to algorithms like Shor’s for factoring, and quantum phase estimation. We first present a lossless local reasoning method for the quantum Fourier transform (QFT) based on Quantum Abstract Interpretation (QAI). Remarkably, for inputs in the computational basis, the concretized abstract output state produced by our analysis exactly matches the true output state. To the best of our knowledge, this is the first approach that demonstrates—within polynomial time—the ability to recover such precise information about the QFT, and, by extension, quantum phase estimation.
Figure 4. Quantum Fourier transform. Swap gates at the end of the circuit that reverse the order of the qubits are not shown.
We use the $\psi=|\psi\rangle\langle\psi|$ for pure state |$\ket{\psi}$,$0.x_{1}x_{2}\cdots x_{n}=\sum_{i=1}^{n}\frac{x_{i}}{2^{i}}$, and the following notation
(26) |$\ket{\psi_{x}}:=\frac{1}{\sqrt{2}}(\ket{0}+e^{2\pi i0.x}\ket{1}).$
@@
In Appendix C, we prove the following using QAI
@@
The postcondition derived from QAI is an abstract state represented as a tuple of density matrices corresponding to pure quantum states. By applying the concretization function from 3.2, we infer that the concrete state lies in the subspace
@@ 6.2. Quantum Phase Estimation
In this section, we present a quantitative local analysis of the Quantum Phase Estimation (QPE) algorithm by combining both backward and forward reasoning techniques. We begin by decomposing the QPE algorithm into three constituent circuits and apply SAQR-QC to reason about each component individually. For an unknown phase $\theta$ and a given constant $k$, SAQR-QC can formally establish that—with probability at least $\frac{4}{\pi^{2}}$—the QPE algorithm produces an output bitstring whose last $k$ bits constitute the optimal $k$-bit approximation to the least significant $k$ bits of any $n$-bit binary representation/approximation of $\theta$.
Figure 5. Quantum Phase Estimation: $U\ket{\psi}=e^{i\theta}\ket{\psi}$. We only consider the circuit without the measurements.
@@
Our analysis focuses on the least significant $k$ qubits, and proceeds in three steps. We will rely on the notation introduced in Equation 36, and apply the QAI-based reasoning framework introduced in § 4.7 to verify the behavior of $\mathbf{C_{1}}$.
Step 1: Reasoning about $\mathbf{C_{2}}$. Let $U$ be the circuit corresponding to $\mathbf{C_{2}}$. For $\mathbf{C_{2}}$, we choose the domain of the predicate as the first $k$ qubits of the first $n$ qubits, together with the last $m$ qubits—the qubits that $U$ acts upon. We use backward reasoning to show
@@
with $\mathscr{A}=(|\tau\rangle\langle\tau|\otimes|\psi\rangle\langle\psi|)$,and |$\ket{\tau}=\ket{\psi_{j_{n}}}\otimes\cdots\otimes\ket{\psi_{j_{n-k+1}\cdots j_{n}}}$. Then,
$M_{\mathscr{A}}=I_{1,\cdots,n-k}\otimes|\tau\rangle\langle\tau|\otimes|\psi\rangle\langle\psi|.$
Equation 28 is equivalent to
To see this, we observe
Let $V=H(1)^{-1}CR_{2}(2,1)^{-1}CR_{3}(3,1)^{-1}\cdots H(n-k+2)^{-1}$ means
$U_{k}=H(n-k+1)^{-1}\cdots H(n-1)^{-1}CR_{2}(n,n-1)^{-1}H(n)^{-1},$
Then
This argument proves Equation 28.
Remark:
Matrix multiplication will be efficient for constant dimension. Here, the dimension of matrices is $2^{k}$, which is a constant independent of $n$.
Step 2: Reasoning about SWAP gates. Now the precondition of Equation 28, $\{\mathscr{A}|\mathcal{I}\}$, becomes the postconditon of the SWAP gates. Since the action of the SWAP gates only changes the last $k$ qubits into the first $k$ qubits, we may just write the precondition of the SWAP gates as $\{\mathscr{A}^{\prime}|\mathcal{I}\}$ with $\mathscr{A}^{\prime}=(|\tau\rangle\langle\tau|\otimes|\psi\rangle\langle\psi|)$ on the first $k$ qubits and the last $m$ qubits. These arguments prove the following
@@
where we use $\mathbf{S}$ to denote the SWAP gates.666 We change the abstract domain for simplicity of presentation. This change will not affect our statement’s correctness because the SWAP gates’ action is clear. If we want to fix the abstract domain, we can consider $(s_{1},\cdots,s_{m})$ with $s_{1}$ being the last $k$ qubits together with the last $m$ qubits; $s_{2}$ being the result of applying the first SWAP gate on $s_{1}$; $\cdots$; $s_{m}$ being the result of applying the last SWAP gate on $s_{m-1}$.
Step 3: Reasoning about $\mathbf{C_{1}}$. For $\mathbf{C_{1}}$, we choose the domain of the predicate as the first $k$ qubits of the first $n$ qubits together with the last $m$ qubits—the qubits on which $U$ applied. We will use our proof rules to show
@@
$r$ $=\Pi_{t=1}^{k}\cos^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]$
@@
|$\ket{\omega}$ $=\frac{1}{2^{k/2}}(\ket{0}+e^{2\pi i2^{n-1}\theta}\ket{1})\otimes\cdots\otimes(\ket{0}+e^{2\pi i2^{n-k}\theta}\ket{1}).$
We first use quantum abstract interpretation (YP21, ) to prove
After applying the first $H$ gates, the post-condition becomes
Similarly for $CU^{2^{n-k+1}}\cdots CU^{2^{0}}$. This argument proves Equation 31.
To prove Equation 30, we observe
and
where
$V(C_{1})$ $=CU^{2^{n-k}}\cdots CU^{2^{n-1}}$
$x$ $=\langle 0\cdots 0\psi|V(C_{1})^{{\dagger}}|\tau\psi\rangle\langle\tau\psi|\ket{0\cdots 0\psi}$
$=|\langle\tau|\omega\rangle|^{2}$
$=\Pi_{t=1}^{k}|\frac{(\langle 0|+e^{-2\pi i0.j_{n-t+1}\cdots j_{n}}\langle 1|)(\ket{0}+e^{2\pi i2^{n-t}\theta}\ket{1}}{2}|^{2}$
$=\Pi_{t=1}^{k}\cos^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]$
$=r.$
$r$ $=\Pi_{t=1}^{k}\cos^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]=\Pi_{t=1}^{k}\frac{\sin^{2}[2(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]}{4\sin^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]}$
$=\Pi_{t=1}^{k}\frac{\sin^{2}[2^{n-t+1}\theta-0.j_{n-t+2}\cdots j_{n})\pi]}{4\sin^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]}=\frac{\sin^{2}(2^{n}\theta\pi)}{4^{k}\sin^{2}[(2^{n-k}\theta-0.j_{n-k+1}\cdots j_{n})\pi]}.$
Let $U\ket{\psi}=e^{i\theta}\ket{\psi}$, and $\theta=\frac{a}{2^{n}}+\epsilon$ with $-\frac{1}{2^{n+1}}\leq\epsilon\leq\frac{1}{2^{n+1}}$ and $a$ is an integer with binary representation $a_{1}a_{2}\cdots a_{n}$. For $j_{n-k}\cdots j_{n}=a_{n-k}\cdots a_{n}$, we have
@@
Together with the Seq Rule and the Con Rule, the proved result can be interpreted as follows: for any constant $k$, the last $k$ bits of the output will—with probability at least $\frac{4}{\pi^{2}}$—match the best $k$-bit binary approximation to the least significant $k$ bits of the phase $\theta$, provided that the first $k$ input qubits are initialized to |$\ket{0,\ldots,0}$, regardless of the state of the remaining $n-k$ qubits.

### v3

@@ 4. Quantitative Reasoning of general GHZ circuit
We analyze a generalized GHZ circuit where half the gates are arbitrary single-qubit unitaries, producing a highly entangled, densely parameterized state beyond the scope of classical simulation methods such as the Gottesman–Knill theorem (Aaronson_2004, ). Using the scalable techniques developed herein, our analysis proceeds by first applying the “warm-up” method from §˜3.6 (Equation˜13), which enables compositional and projection-free reasoning over subsystems, thereby supporting tractable approximation of the circuit’s complex output state.
Figure 3. GHZ circuit with unitaries $U_{i}$.
In the second phase, we leverage Quantum Abstract Interpretation (QAI), employing the Assert rule from ˜2.2 to confine the output state within a two-dimensional subspace. This critical dimensionality reduction facilitates precise amplitude approximations within our quantitative reasoning framework. Strikingly, the resulting characterization matches the exact wavefunction up to a phase, underscoring that reasoning in SAQR-QC is not merely procedural but requires judicious predicate selection aligned with the circuit’s structure—demonstrating both precision and flexibility when applied with insight.
We select the domain $(\{1,2\},\{2,3\},\cdots,\{n-1,n\})$ for reasoning about Figure 3.
STEP 1. We first select the precondition to be $\{\mathscr{A}|\mathcal{P}\}$
@@
After the first $H$ gate, we can use inequality 12 to derive the following:
$\sum_{i}{A_{s_{i}}}\otimes I_{[n]\setminus\{i,i+1\}}\leq H^{{\dagger}}\sum_{i}{B_{i,i+1}}\otimes I_{[n]\setminus\{i,i+1\}}H$
@@
The final inequality above matches inequality 14 on a term-by-term basis. We can satisfy inequality 14 when each of the following single-term inequalities hold:
@@
Each inequality can be realized as an equality, yielding the postcondition $\mathscr{B}\mid\mathcal{Q}$},
@@
At this point, we can satisfy inequality 14 by satisfying the following smaller inequalities:
@@
We can choose $C_{i,i+1}$ to be $|++\rangle\langle++|$ for $i>2$, and for remaining two-term-per-side inequality, we can derive the following:
@@
(coinciding with the pre-state predicate $\mathscr{B}=(B_{1,2},B_{2,3},\dots,B_{n-1,n})$), and the post-state predicate is
@@
Equation˜18 illustrates an advantage of our choice of predicates. Because of properties such as those given in Equation˜17, Equation˜12 remains invariant under the application of CNOT gates. This invariance allows us to derive the strongest postcondition, while preserving the local structure of the matrix representation of predicates. As a result, we were able to make choices that made the inequalities that we worked with tight (or saturated, i.e., satified as equalities), making it easier to determine the postcondition. The right-hand side of Equation˜18 continues to serve as the predicate of local observables, as reasoning continues about the remaining CNOT gates.
After applying the CNOT gate on $q_{1}$ and $q_{r}$, the postcondition can be chosen as $(\mathscr{D}\mid\mathcal{S})$, where
$\mathscr{D}$ $=(|0+\rangle\langle 0+|,|++\rangle\langle++|,\cdots,|++\rangle\langle++|)$
@@
[…]
Let the output state be $\rho=|\Psi\rangle\langle\Psi|$. By ˜3.8, our proof of $\mathscr{F}$ implies that
(19) $\frac{n-1}{4}=\sum_{i=1}^{n-1}{\mathrm{Tr}}(|00\rangle\langle 00||++\rangle\langle++|)\leq\sum_{i=1}^{n-1}{\mathrm{Tr}}[\rho_{i,i+1}(\beta_{i}\otimes\beta_{i+1})].$
STEP 2. This step mirrors Step 1, starting from the precondition $((|--\rangle\langle--|,|--\rangle\langle--|,\dots,|--\rangle\langle--|)\mid\mathcal{P})$. The postcondition of the full circuit can then be computed as
[…]
These conditions imply that
@@
${\mathrm{Tr}}[(\psi_{1}\otimes\psi_{2})(\beta_{1}\otimes\beta_{2})]=\frac{1}{2},\ {\mathrm{Tr}}[(\psi_{i}\otimes\psi_{i+1})(\beta_{i}\otimes\beta_{i+1})]=\frac{1}{4}\ \ \forall i>1$
${\mathrm{Tr}}[(\phi_{1}\otimes\phi_{2})(\beta_{1}\otimes\beta_{2})]=0,\ {\mathrm{Tr}}[(\phi_{i}\otimes\phi_{i+1})(\beta_{i}\otimes\beta_{i+1})]=\frac{1}{4}\ \ \forall i>1$
${\mathrm{Tr}}[(\psi_{1}\otimes\psi_{2})(\delta_{1}\otimes\delta_{2})]=0,\ {\mathrm{Tr}}[(\psi_{i}\otimes\psi_{i+1})(\delta_{i}\otimes\delta_{i+1})]=\frac{1}{4}\ \ \forall i>1$
${\mathrm{Tr}}[(\phi_{1}\otimes\phi_{2})(\delta_{1}\otimes\delta_{2})]=\frac{1}{2},\ {\mathrm{Tr}}[(\phi_{i}\otimes\phi_{i+1})(\delta_{i}\otimes\delta_{i+1})]=\frac{1}{4}\ \ \forall i>1.$
Equations˜19 and 20 imply
@@
[…]
Equation˜21 represents a closed-form expression for the circuit’s output with a single unknown real parameter. Such quantitative reasoning cannot be achieved with qualitative methods like QAI alone.
@@ 5. Quantum Phase Estimation
We apply our proof system to analyze Quantum Phase Estimation (QPE), a foundational algorithm central to Shor’s algorithm, whose intricate entanglement and nuanced success bounds have historically resisted scalable formal verification. Our approach unfolds in two stages: first, we establish the correctness of the Quantum Fourier Transform (QFT) via Quantum Abstract Interpretation (QAI); second, we leverage this foundation to analyze the full QPE circuit within our logic framework, deriving a lower bound on its success probability. To our knowledge, this constitutes the first scalable formal verification of QPE and QFT, offering a compositional, tractable methodology for reasoning about both correctness and probabilistic performance in key quantum algorithms.
@@ 5.1. Quantum Fourier transform
The Quantum Fourier Transform (QFT)—the quantum analogue of the discrete Fourier transform (see Figure˜4)—forms the computational core of landmark algorithms such as Shor’s factoring and quantum phase estimation. We introduce a lossless local reasoning framework for the QFT grounded in Quantum Abstract Interpretation (QAI). Our method derives an abstract output state whose concretization exactly coincides with the true QFT output for all computational-basis inputs. To the best of our knowledge, this constitutes the first verification framework that provides proofs polynomial in the size of the program while achieving such exact semantic correspondence for the QFT—and, by extension, for quantum phase estimation.
Figure 4. Quantum Fourier Transform: swap gates that reverse the qubit order at the circuit’s end are omitted.
The circuit uses Hadamard and phase gates $R_{m}$; for $m>2$, $R_{m}$ lies outside the Clifford group. See Subsection 2.4 for definitions of $H$ and $R_{m}$.
[…]
@@
In Appendix E, we prove the following using QAI
@@
The postcondition derived from QAI is an abstract state represented as a tuple of density matrices corresponding to pure quantum states. By applying the concretization function from ˜2.2, we infer that the concrete state lies in the subspace
@@ 5.2. Quantum Phase Estimation
In this section, we present a quantitative local analysis of the Quantum Phase Estimation (QPE) algorithm by combining both backward and forward reasoning techniques. We begin by decomposing the QPE algorithm into three constituent circuits and apply SAQR-QC to reason about each component individually—see Figure˜5. For an unknown phase $\theta$ and a given constant $k$, SAQR-QC can formally establish that—with probability at least $\frac{4}{\pi^{2}}$—the QPE algorithm produces an output bitstring whose last $k$ bits constitute the optimal $k$-bit approximation to the least significant $k$ bits of any $n$-bit binary representation/approximation of $\theta$.
[…]
@@
Our analysis focuses on the least significant $k$ qubits, and proceeds in three steps. We will rely on the notation introduced in Equation˜30, and apply the QAI-based reasoning framework introduced in §˜3.6 to verify the behavior of $\mathbf{C_{1}}$.
Step 1: Reasoning about $\mathbf{C_{2}}$. Let $U$ be the circuit corresponding to $\mathbf{C_{2}}$. For $\mathbf{C_{2}}$, we choose the domain of the predicate as the first $k$ qubits of the first $n$ qubits, together with the last $m$ qubits—the qubits that $U$ acts upon. In Appendix F.1, we use backward reasoning to show
@@
[…]
Remark: Matrix multiplication will be efficient for constant dimension. Here, the dimension of matrices is $2^{k}$, which is a constant independent of $n$.
Step 2: Reasoning about SWAP gates. Now the precondition of Equation˜24, $\{\mathscr{A}|\mathcal{I}\}$, becomes the postconditon of the SWAP gates. Since the action of the SWAP gates only changes the last $k$ qubits into the first $k$ qubits, we may just write the precondition of the SWAP gates as $\{\mathscr{A}^{\prime}|\mathcal{I}\}$ with $\mathscr{A}^{\prime}=(|\tau\rangle\langle\tau|\otimes|\psi\rangle\langle\psi|)$ on the first $k$ qubits and the last $m$ qubits. These arguments prove
@@
where we use $\mathbf{S}$ to denote the SWAP gates. [^3] We change the abstract domain for simplicity of presentation. This change will not affect our statement’s correctness because the SWAP gates’ action is clear. If we want to fix the abstract domain, we can consider $(s_{1},\cdots,s_{m})$ with $s_{1}$ being the last $k$ qubits together with the last $m$ qubits; $s_{2}$ being the result of applying the first SWAP gate on $s_{1}$; $\cdots$; $s_{m}$ being the result of applying the last SWAP gate on $s_{m-1}$.
Step 3: Reasoning about $\mathbf{C_{1}}$. For $\mathbf{C_{1}}$, we select the predicate domain to include the first $k$ qubits of the initial $n$ and the last $m$ qubits where $U$ acts. In Appendix F.2, we apply our proof rules to demonstrate
@@
$r$ $=\Pi_{t=1}^{k}\cos^{2}[(2^{n-t}\theta-0.j_{n-t+1}\cdots j_{n})\pi]=\frac{\sin^{2}(2^{n}\theta\pi)}{4^{k}\sin^{2}[(2^{n-k}\theta-0.j_{n-k+1}\cdots j_{n})\pi]}$
@@
[…]
@@
[…]